The area in square meters for a triangular sail is given by A(x)=x^2+5x+6
Find A(5)? If the height of the sail is x+3 meters then what is the length of the base of the sail?

Answer by edjones(8007)

A(x)=x^2+5x+6
A(5)=5^2+5*5+6
=25+25+6
=56
.
A=.5ab
.5(5+3)b=56
4b=56
b=14 m
.
Ed